# Distributed Computing From First Principles

### v1

@@ Acknowledgement
We have provided credits for help received in writing this book.
I want to thank Anselm Eickhoff and Charles Krempeaux, who were responsive to my questions about CRDT.
I would also like to thank Harrington Joseph, who suggested that I focus on gaining a low-level understanding of parallel programming.
@@
Have you ever wondered how a typical distributed system works under the hood? Are you looking for a pedagogical guide with complete implementations and tricks of the trade? Look no further and read my writing on the topic. We have implemented several foundational algorithms in Distributed Computing. This paradigm has become ubiquitous in the industry where multiple systems interact to solve computational tasks, with applications including distributed stores (databases), IoT sensor networks, and the Internet. The exponential growth in deploying Distributed systems for solving real-world problems has led to a resurgence in understanding Distributed Computing from scratch 111 My research in Distributed Systems is a two-phased process: The first phase was covered in the https://kenluck2001.github.io/blog_post/distributed_computing_from_first_principles.html. The second phase is the current book that you are reading. A multi-phased approach to research has helped to minimize risk.. Consensus is a foundational requirement for distributed systems, allowing disparate nodes to synthesize a consistent global perspective from their local operations. This collaborative agreement forms the very objective of distributed algorithms.
@@ 1.1 Algorithm Implementations
Some of my implementations are in the playground folder, where you can follow my thought process during incremental development. Some of the implementations in the playground contain faulty solutions due to incorrect assumptions. However, the implementations referenced directly in this book are verified by the author and relatively free of obvious errors. We settled on OpenMPI as the underlying library to provide low-level functionality. This choice is preferable to using a Remote Procedural Call, which obscures the message-passing paradigm by passing arguments as messages. Although conceptually similar, MPI provides a structured way of implementing our Distributed algorithms. Hence, all implementations will strictly use OpenMPI in C. Event-based programming based on message passing serves as a powerful abstraction for building Distributed systems. Therefore, we are using this approach for all of our implementations. For the sake of pedagogy, we have omitted tracing and invariant proofs. While these are important in Distributed systems scholarship, we anticipate that users will learn by doing and not be hindered by excessive mathematics. However, for those who wish to focus more on writing proofs, we recommend this resource [33].
@@ 2.1 Network Programming
The socket is the fundamental file descriptor for networking. Recalling the Unix philosophy that everything is a file, the ‘socket()‘ function returns a descriptor. Sending a message over a network becomes analogous to writing to a file stream, while receiving a message is comparable to reading a file stream. However, in this context, the network acts as the channel for this communication. There are two basic socket types, as outlined by [27]:
@@
When communicating between nodes in Distributed systems, a channel is necessary to provide a medium for message transfer. The types of channels in use in Distributed systems [37] include:
@@ 2.2 Parallel Programming
In Unix, ‘fork()‘ creates a new process, where a parent process creates a new child process. The parent process must wait for the child process to exit, allowing the child process to terminate properly. This behaviour can become problematic if the child process becomes defunct (a zombie) and the parent process ignores waiting. In some systems, the ‘init‘ process reaps (destroys) defunct processes. A child process becomes a zombie until the parent process waits or the parent ignores the ‘SIGCHLD‘ signal [27, 28]. Ignoring the waiting for the child process to exit in the parent process:
@@
Processes and threads are core foundations for building Distributed computing systems [9, 12]. Processes are units of work distribution. Threads are units of concurrency. Parallel programming presents significant challenges, including data sharing, coordination, deadlock, lock granularity, and others [9, 12]. It is possible to have multiple threads within a single process. Some benefits of this include [9, 12]:
@@
Another combination is to have multiple processes within a node. Some benefits of this include [9, 12]:
@@
There are different forms of parallelism (task, functional, loop, data-flow) [9]. Java’s popular fork-join framework is based on the divide and conquer paradigm, which is useful, where the problem can be decomposed in this sub-problem structure [9].
@@
Guaranteeing reproducibility can be a desirable goal. Quasi-randomness may be acceptable in some contexts. A parallel program can exhibit the following characteristics to address this behavior [9]:
@@
Ideal parallelism ($ip$) is given by: $ip=work/span$ For a sequential program, the ideal parallelism is 1. This requires having a computation graph [9]. Unfortunately, we don’t always have a computation graph. In such cases, we can use Amdahl’s law, which states that the speedup of a threaded program is limited by the sequential portion of the computation across all processes [9, 12]. There are also alternatives like share-nothing architecture (using the message-passing paradigm), e.g., actor model, and distributed actors, which make use of message passing. Non-blocking approaches using ‘getAndAdd()‘ and ‘compareAndSet()‘ are also available [9].
Properties of a Parallel Program [9]
@@ 2.2.1 Process Communication
Synchronization guarantees the order of access of multiple threads to a shared resource. Synchronization outside a computer can be verified in real life by a clock. However, achieving computer synchronization without a clock is challenging, as there are also issues with the accuracy of clocks due to drifts. It is possible to do so using computational graphs, especially in a parallel program. While it is easy to know the order of execution within a process, it is challenging to ascertain the order of execution in a threaded program. Synchronization is achievable by signaling or instructing the other party to wait until a condition gets met. There are traditional synchronization problems posed in the literature with wide-applicability. These include the producer-consumer problem (with multiple variants), the reader-writer problem, the starvation problem (bounded wait on a semaphore), and the dining philosophers problem [9, 23, 28].
A barrier is a construct that blocks multiple threads and releases every blocked thread upon the arrival of the last one [23]. This paradigm is commonly known as a phaser in Java and is widely employed in MPI for creating collective operations. This construct can also be used for pipelining [9, 23, 28]. A good book on interprocess communication is Beej’s Guide to Unix IPC by Brian Hall. There are multiple venues for sharing information between processes in Unix.
@@ 2.2.2 OpenMPI
OpenMPI [6, 7, 8, 10] is an implementation of the message-passing paradigm for the exchange of information between processes, widely used in rendering farms, high-performance computing, and Scientific computing. For example, Curie, a French supercomputer, makes extensive use of OpenMPI for processing workflow. One of the best resources for learning MPI on the Internet can be found on OpenMPI website.
The number of processes, $size$, is determined during the setup of the OpenMPI communicator. However, adjustments can be made using sophisticated process group management. Every process is assigned a rank, ranging from 0 to $size$-1.
@@
Unlike the two-sided communication model, which requires a sender and a receiver for data transfer, the one-sided communication model allows a process to directly access another process’s memory space. Consequently, only one process initiates communication directly, minimizing data transfer and CPU intervention. However, there is a caveat that the ordering of RMA operations cannot be safely guaranteed [6, 10]. In our implementation of Distributed shared primitives, we made extensive use of one-sided communication. While RMA is inherently non-blocking, we are currently employing it in a blocking mode using locks. Implementing an epoch with MPI_Win_fence might be a more suitable approach in our logic to fully leverage the non-blocking features. MPI_Compare_and_swap could also present a better alternative [6, 10]. If one-sided communication is used in place of the default two-sided communication, the concept of explicit acknowledgments (ACKS or receipt confirmations) as in two-sided communication doesn’t directly apply. The communication is more akin to direct memory manipulation. Here is an example of an epoch:
@@
Distributed Systems [9, 12] are sets of nodes (devices) connected by communication links that operate as a single, cohesive system. Although each device functions locally and independently, the overall system appears as a unified global entity, providing the ”single view illusion.” Examples include the Internet, edge computing environments, mobile networks, and sensor networks, as illustrated in Figure 3.1.
@@
Consensus is achieved when every node agrees on the same value. Furthermore, the decided value must be one of the values that was initially proposed. Consensus is equivalent to atomic commits (atomic broadcast) [4].
Reasoning about Distributed computation can be inherently complex, necessitating the use of simplifying models. One such model involves specifying the failure assumptions, which can encompass various modes of process failure [4]:
@@
NoSQL database systems tend to be more readily adaptable to distributed concepts, such as key-value stores and document stores [4, 12]. Several database architectures can ensure data availability across nodes through replication strategies (e.g., master-slave, master-master, buddy replication).
A replicated state machine is a technique used to guarantee availability in the face of failures by mirroring data and state across multiple nodes. It adheres to the following fundamental properties [4]:
@@
Nodes within a Distributed system can assume various roles [37]:
@@ 3.1.1 FLP Impossibility of Consensus
The Fischer-Lynch-Paterson (FLP) impossibility result describes the inherent limitations of achieving consensus in asynchronous distributed systems. It states that in an asynchronous system subject to even a single crash-stop failure, it is impossible to guarantee that a consensus algorithm will always satisfy the following three crucial properties [4, 12, 37]:
@@
The FLP result highlights the fundamental challenges of achieving reliable consensus in systems with unpredictable timing, applicable to asynchronous, synchronous, and partially synchronous systems [4, 12, 37]:
@@
  • In a synchronous system, consensus cannot be solved if $N-1$ nodes fail (where $N$ is the total number of nodes).
@@ 3.1.2 Two Generals Problem
The Two Generals Problem, also known as the Coordinated Attack Problem, illustrates the difficulties of achieving agreement over an unreliable communication channel. Two generals must agree on a time to launch a joint attack, communicating only through messengers who might be captured or delayed [4, 12, 37]. Mapping this problem to a real-world Distributed system reveals the following analogous challenges [4, 12, 37]:
@@
The Two Generals Problem demonstrates that it is impossible to guarantee that two parties can reach a perfect consensus in an asynchronous system when there is even a possibility of a single faulty communication link [4, 12, 37].
@@ 3.2 Logical Clocks
If event ’a’ happens-before event ’b’ on the same process, we denote this as $a\rightarrow b$, where ’a’ could be a send event and ’b’ a subsequent deliver event. This ”happens-before” relation is fundamental for inferring event order in the absence of a global clock [4]. Two primary types of logical clocks are commonly used:
@@ 3.2.1 Lamport clock
The Lamport clock utilizes the ”happens-before” relation to establish a total order of events. Timestamps generated by a Lamport clock have a defined less-than relationship ($<$), allowing for the ordering of any two events. However, because the timestamp is a single integer, it does not capture information about non-causality (concurrency).
@@
  • Recv(m) (LP=$\max(0+1,3)=3$): Upon receiving message $m$ with a timestamp of 3 from P1, Process P2 updates its Lamport clock to the maximum of its local clock incremented by one (1) and the received timestamp (3), resulting in a timestamp of 3 for the received event.
@@
  • Recv(n) (LP=$\max(1+1,5)=5$): Upon receiving message $n$ with a timestamp of 5 from P2, Process P3 updates its Lamport clock to the maximum of its local clock incremented by one (2) and the received timestamp (5), resulting in a timestamp of 5 for the received event.
@@ 3.2.2 Vector clock
The Vector clock, in contrast, is based on partial order and effectively captures both causality and non-causality [13]. The ability to detect non-causality is crucial for identifying events that occur concurrently.
@@ 3.3 Failure Detector
A failure detector is a mechanism designed to identify faults (specifically, process crashes) in asynchronous distributed systems. The introduction of failure detectors leads to a refined classification of asynchronous systems known as ”Timed Asynchronous systems” or ”Partially Synchronous systems.” These systems exhibit the following characteristics [4]:
@@
A typical failure detection algorithm operates as follows [4, 12, 37]:
@@
In asynchronous systems, consensus and atomic broadcast become solvable with the aid of a failure detector. For a failure detector to be practically useful, it must satisfy certain requirements with varying degrees of certainty [4, 12, 37]:
@@ 3.4 Graceful Degradation
Failure tolerance is at the core of building fault-handling mechanisms using resiliency patterns (e.g., retry, circuit breaker, fail-fast, bulkhead) [11]. The overarching aim of these patterns is to prevent requests from being directed to faulty resources, thereby enhancing the resilience of the distributed systems.
@@ 4.1 Paxos Algorithm
Paxos is a fundamental algorithm for achieving consensus among multiple nodes across several rounds of communication. It provides a method for achieving abortable consensus in a distributed system with the following key properties [12]:
@@
The algorithm generally follows these steps [4, 37]:
@@
The following concurrency-related properties [4, 12, 37] are generally applicable in parallel and concurrent programming paradigms:
@@
The main characteristics of the Paxos algorithm are [4, 37, 12]:
@@
The Single Valued Paxos algorithm proceeds as follows [4, 12]:
@@
Properties of the Paxos algorithm [4, 12]:
@@
Figure 4.1: Initialization of states at the internal start of Paxos Algorithm [4].
@@
Figure 4.2: Sequence of message flow in a Paxos Algorithm [4].
It’s worth noting that there are several variations of Single Value Paxos, including Fast Paxos, Egalitarian Paxos, and Flexible Paxos [37], each offering different performance characteristics or fault-tolerance trade-offs.
@@
The Sequence Paxos algorithm extends the Single Value Paxos to achieve agreement on a sequence of values (often representing a log of events) in a consistent order across all nodes [4, 37, 12]. The general approach involves:
@@
Figure 4.3: Initialization of internal states at the start of Sequence Paxos Algorithm [4].
@@
Figure 4.4: Message flow in a Sequence Paxos Algorithm [4].
@@
Figure 4.5: Maintaining prefix of log invariants in a Sequence Paxos Algorithm [4].

### v2

@@ Acknowledgement
Archit Goyal reached out and provided source code on CRDT among other major reviewing and coding contributions. Archit is a Senior Software Engineer at LinkedIn, deeply involved in building and productionizing large-scale distributed data systems. He is one of the early engineers behind LinkedIn’s Flink Batch platform, powering Ads AI workloads and high-throughput, low-latency pipelines. Much of his work focuses on scaling stateful computations, disaggregated shuffle, checkpointing trade-offs, failure recovery, resource fairness, and multi-cluster observability.
I want to thank Anselm Eickhoff and Charles Krempeaux, who were responsive to my questions about CRDT. Also, I would also like to thank Harrington Joseph, who suggested that I focus on gaining a low-level understanding of parallel programming. We provided credits for both anonymous reviewers that provided tangible support in making this manuscript.
@@
Have you ever wondered how a typical distributed system works under the hood? Are you looking for a pedagogical guide with complete implementations and tricks of the trade? Look no further and read my writing on the topic. We have implemented several foundational algorithms in Distributed Computing. This paradigm has become ubiquitous in the industry where multiple systems interact to solve computational tasks, with applications including distributed stores (databases), IoT sensor networks, and the Internet. The exponential growth in deploying Distributed systems for solving real-world problems has led to a resurgence in understanding Distributed Computing from scratch  [^1] My research in Distributed Systems is a two-phased process: The first phase was covered in the https://kenluck2001.github.io/blog_post/distributed_computing_from_first_principles.html. The second phase is the current book that you are reading. A multi-phased approach to research has helped to minimize risk.. Consensus is a foundational requirement for distributed systems, allowing disparate nodes to synthesize a consistent global perspective from their local operations. This collaborative agreement forms the very objective of distributed algorithms.
@@ 1.1 Algorithm Implementations
Some of my implementations are in the playground folder, where you can follow my thought process during incremental development. Some of the implementations in the playground contain faulty solutions due to incorrect assumptions. However, the implementations referenced directly in this book are verified by the author and relatively free of obvious errors. We settled on OpenMPI as the underlying library to provide low-level functionality. This choice is preferable to using a Remote Procedural Call, which obscures the message-passing paradigm by passing arguments as messages. Although conceptually similar, MPI provides a structured way of implementing our Distributed algorithms. Hence, all implementations will strictly use OpenMPI in C. Event-based programming based on message passing serves as a powerful abstraction for building Distributed systems. Therefore, we are using this approach for all of our implementations. For the sake of pedagogy, we have omitted tracing and invariant proofs. While these are important in Distributed systems scholarship, we anticipate that users will learn by doing and not be hindered by excessive mathematics. However, for those who wish to focus more on writing proofs, we recommend this resource [lynch1993].
@@ 2.1 Network Programming
The socket is the fundamental file descriptor for networking. Recalling the Unix philosophy that everything is a file, the ‘socket()‘ function returns a descriptor. Sending a message over a network becomes analogous to writing to a file stream, while receiving a message is comparable to reading a file stream. However, in this context, the network acts as the channel for this communication. There are two basic socket types, as outlined by [brian2020]:
@@
When communicating between nodes in Distributed systems, a channel is necessary to provide a medium for message transfer. The types of channels in use in Distributed systems [petrov2019] include:
@@ 2.2 Parallel Programming
In Unix, ‘fork()‘ creates a new process, where a parent process creates a new child process. The parent process must wait for the child process to exit, allowing the child process to terminate properly. This behaviour can become problematic if the child process becomes defunct (a zombie) and the parent process ignores waiting. In some systems, the ‘init‘ process reaps (destroys) defunct processes. A child process becomes a zombie until the parent process waits or the parent ignores the ‘SIGCHLD‘ signal [brian2020, brian2015]. Ignoring the waiting for the child process to exit in the parent process:
@@
Processes and threads are core foundations for building Distributed computing systems [pcdp2020, alinush2015]. Processes are units of work distribution. Threads are units of concurrency. Parallel programming presents significant challenges, including data sharing, coordination, deadlock, lock granularity, and others [pcdp2020, alinush2015]. It is possible to have multiple threads within a single process. Some benefits of this include [pcdp2020, alinush2015]:
@@
Another combination is to have multiple processes within a node. Some benefits of this include [pcdp2020, alinush2015]:
@@
There are different forms of parallelism (task, functional, loop, data-flow) [pcdp2020]. Java’s popular fork-join framework is based on the divide and conquer paradigm, which is useful, where the problem can be decomposed in this sub-problem structure [pcdp2020].
@@
Guaranteeing reproducibility can be a desirable goal. Quasi-randomness may be acceptable in some contexts. A parallel program can exhibit the following characteristics to address this behavior [pcdp2020]:
@@
Ideal parallelism ($ip$) is given by: $ip=work/span$ For a sequential program, the ideal parallelism is 1. This requires having a computation graph [pcdp2020]. Unfortunately, we don’t always have a computation graph. In such cases, we can use Amdahl’s law, which states that the speedup of a threaded program is limited by the sequential portion of the computation across all processes [pcdp2020, alinush2015]. There are also alternatives like share-nothing architecture (using the message-passing paradigm), e.g., actor model, and distributed actors, which make use of message passing. Non-blocking approaches using ‘getAndAdd()‘ and ‘compareAndSet()‘ are also available [pcdp2020].
Properties of a Parallel Program [pcdp2020]
@@ 2.2.1 Process Communication
Synchronization guarantees the order of access of multiple threads to a shared resource. Synchronization outside a computer can be verified in real life by a clock. However, achieving computer synchronization without a clock is challenging, as there are also issues with the accuracy of clocks due to drifts. It is possible to do so using computational graphs, especially in a parallel program. While it is easy to know the order of execution within a process, it is challenging to ascertain the order of execution in a threaded program. Synchronization is achievable by signaling or instructing the other party to wait until a condition gets met. There are traditional synchronization problems posed in the literature with wide-applicability. These include the producer-consumer problem (with multiple variants), the reader-writer problem, the starvation problem (bounded wait on a semaphore), and the dining philosophers problem [pcdp2020, downey2016, brian2015].
A barrier is a construct that blocks multiple threads and releases every blocked thread upon the arrival of the last one [downey2016]. This paradigm is commonly known as a phaser in Java and is widely employed in MPI for creating collective operations. This construct can also be used for pipelining [pcdp2020, downey2016, brian2015]. A good book on interprocess communication is Beej’s Guide to Unix IPC by Brian Hall. There are multiple venues for sharing information between processes in Unix.
@@ 2.2.2 OpenMPI
OpenMPI [cs598-s152022, llnlgov2022, mpitutorial2022, eijkhout2022] is an implementation of the message-passing paradigm for the exchange of information between processes, widely used in rendering farms, high-performance computing, and Scientific computing. For example, Curie, a French supercomputer, makes extensive use of OpenMPI for processing workflow. One of the best resources for learning MPI on the Internet can be found on OpenMPI website.
The number of processes, sizesize, is determined during the setup of the OpenMPI communicator. However, adjustments can be made using sophisticated process group management. Every process is assigned a rank, ranging from 0 to sizesize-1.
@@
Unlike the two-sided communication model, which requires a sender and a receiver for data transfer, the one-sided communication model allows a process to directly access another process’s memory space. Consequently, only one process initiates communication directly, minimizing data transfer and CPU intervention. However, there is a caveat that the ordering of RMA operations cannot be safely guaranteed [cs598-s152022, eijkhout2022]. In our implementation of Distributed shared primitives, we made extensive use of one-sided communication. While RMA is inherently non-blocking, we are currently employing it in a blocking mode using locks. Implementing an epoch with MPI_Win_fence might be a more suitable approach in our logic to fully leverage the non-blocking features. MPI_Compare_and_swap could also present a better alternative [cs598-s152022, eijkhout2022]. If one-sided communication is used in place of the default two-sided communication, the concept of explicit acknowledgments (ACKS or receipt confirmations) as in two-sided communication doesn’t directly apply. The communication is more akin to direct memory manipulation. Here is an example of an epoch:
@@
Distributed Systems [pcdp2020, alinush2015] are sets of nodes (devices) connected by communication links that operate as a single, cohesive system. Although each device functions locally and independently, the overall system appears as a unified global entity, providing the ”single view illusion.” Examples include the Internet, edge computing environments, mobile networks, and sensor networks, as illustrated in Figure 3.1.
@@
Consensus is achieved when every node agrees on the same value. Furthermore, the decided value must be one of the values that was initially proposed. Consensus is equivalent to atomic commits (atomic broadcast) [kthcourseds].
Reasoning about Distributed computation can be inherently complex, necessitating the use of simplifying models. One such model involves specifying the failure assumptions, which can encompass various modes of process failure [kthcourseds]:
@@
NoSQL database systems tend to be more readily adaptable to distributed concepts, such as key-value stores and document stores [kthcourseds, alinush2015]. Several database architectures can ensure data availability across nodes through replication strategies (e.g., master-slave, master-master, buddy replication).
A replicated state machine is a technique used to guarantee availability in the face of failures by mirroring data and state across multiple nodes. It adheres to the following fundamental properties [kthcourseds]:
@@
Nodes within a Distributed system can assume various roles [petrov2019]:
@@ 3.1.1 FLP Impossibility of Consensus
The Fischer-Lynch-Paterson (FLP) impossibility result describes the inherent limitations of achieving consensus in asynchronous distributed systems. It states that in an asynchronous system subject to even a single crash-stop failure, it is impossible to guarantee that a consensus algorithm will always satisfy the following three crucial properties [kthcourseds, alinush2015, petrov2019]:
@@
The FLP result highlights the fundamental challenges of achieving reliable consensus in systems with unpredictable timing, applicable to asynchronous, synchronous, and partially synchronous systems [kthcourseds, alinush2015, petrov2019]:
@@
  • In a synchronous system, consensus cannot be solved if N1N-1 nodes fail (where $N$ is the total number of nodes).
@@ 3.1.2 Two Generals Problem
The Two Generals Problem, also known as the Coordinated Attack Problem, illustrates the difficulties of achieving agreement over an unreliable communication channel. Two generals must agree on a time to launch a joint attack, communicating only through messengers who might be captured or delayed [kthcourseds, alinush2015, petrov2019]. Mapping this problem to a real-world Distributed system reveals the following analogous challenges [kthcourseds, alinush2015, petrov2019]:
@@
The Two Generals Problem demonstrates that it is impossible to guarantee that two parties can reach a perfect consensus in an asynchronous system when there is even a possibility of a single faulty communication link [kthcourseds, alinush2015, petrov2019].
@@ 3.2 Logical Clocks
If event ’a’ happens-before event ’b’ on the same process, we denote this as $a\rightarrow b$, where ’a’ could be a send event and ’b’ a subsequent deliver event. This ”happens-before” relation is fundamental for inferring event order in the absence of a global clock [kthcourseds]. Two primary types of logical clocks are commonly used:
@@ 3.2.1 Lamport clock
The Lamport clock utilizes the ”happens-before” relation to establish a total order of events. Timestamps generated by a Lamport clock have a defined less-than relationship (<<), allowing for the ordering of any two events. However, because the timestamp is a single integer, it does not capture information about non-causality (concurrency).
@@
[…]
@@
[…]
@@ 3.2.2 Vector clock
The Vector clock, in contrast, is based on partial order and effectively captures both causality and non-causality [attiya2004]. The ability to detect non-causality is crucial for identifying events that occur concurrently.
@@ 3.3 Failure Detector
A failure detector is a mechanism designed to identify faults (specifically, process crashes) in asynchronous distributed systems. The introduction of failure detectors leads to a refined classification of asynchronous systems known as ”Timed Asynchronous systems” or ”Partially Synchronous systems.” These systems exhibit the following characteristics [kthcourseds]:
@@
A typical failure detection algorithm operates as follows [kthcourseds, alinush2015, petrov2019]:
@@
In asynchronous systems, consensus and atomic broadcast become solvable with the aid of a failure detector. For a failure detector to be practically useful, it must satisfy certain requirements with varying degrees of certainty [kthcourseds, alinush2015, petrov2019]:
@@ 3.4 Graceful Degradation
Failure tolerance is at the core of building fault-handling mechanisms using resiliency patterns (e.g., retry, circuit breaker, fail-fast, bulkhead) [microazure]. The overarching aim of these patterns is to prevent requests from being directed to faulty resources, thereby enhancing the resilience of the distributed systems.
@@ 4.1 Paxos Algorithm
Paxos is a fundamental algorithm for achieving consensus among multiple nodes across several rounds of communication. It provides a method for achieving abortable consensus in a distributed system with the following key properties [alinush2015]:
@@
The algorithm generally follows these steps [kthcourseds, petrov2019]:
@@
The following concurrency-related properties [kthcourseds, alinush2015, petrov2019] are generally applicable in parallel and concurrent programming paradigms:
@@
The main characteristics of the Paxos algorithm are [kthcourseds, petrov2019, alinush2015]:
@@
The Single Valued Paxos algorithm proceeds as follows [kthcourseds, alinush2015]:
@@
Properties of the Paxos algorithm [kthcourseds, alinush2015]:
@@
Figure 4.1: Initialization of states at the internal start of Paxos Algorithm [kthcourseds].
@@
Figure 4.2: Sequence of message flow in a Paxos Algorithm [kthcourseds].
It’s worth noting that there are several variations of Single Value Paxos, including Fast Paxos, Egalitarian Paxos, and Flexible Paxos [petrov2019], each offering different performance characteristics or fault-tolerance trade-offs.
@@
The Sequence Paxos algorithm extends the Single Value Paxos to achieve agreement on a sequence of values (often representing a log of events) in a consistent order across all nodes [kthcourseds, petrov2019, alinush2015]. The general approach involves:
@@
Figure 4.3: Initialization of internal states at the start of Sequence Paxos Algorithm [kthcourseds].
@@
Figure 4.4: Message flow in a Sequence Paxos Algorithm [kthcourseds].
@@
Figure 4.5: Maintaining prefix of log invariants in a Sequence Paxos Algorithm [kthcourseds].